OpenStudy (anonymous):

what is the equation in standard form? (4x)^2+y^2-48x-4y+48=0

OpenStudy (anonymous):

\[4x^2+y^2-48x-4y+48=0\] is an ellipse you want the standard form of this?

OpenStudy (anonymous):

\[4x^2-48x+y^2-4y=-48\] \[4(x^2-12x)+y^2-4y=-48\] \[4(x^2-12x+36)+y^2-4y+4=-48+144+4=100\] \[4(x-6)^2+(y-2)^2=100\]

OpenStudy (anonymous):

divide by 100 to get \[\frac{(x-6)^2}{25}+\frac{(y-2)^2}{100}=1\]

OpenStudy (anonymous):

thank you!!!